OpenStudy (anonymous):

explicit rule for -8, -2, 4, 10, 16

OpenStudy (calculusxy):

add 6

OpenStudy (anonymous):

Differences: 6, 6, 6, 6 -> Linear y = mx + b Need m, b -8 = m(1)+b 2 = m(2)+b -8 = m + b 2 = 2m + b Now, solve for m and b. m = -8 - b 2 = 2(-8-b)+b 2 = -16 - 2b + b = -16 - b 2 = -16 - b 14 = -b b = -14 m = -8 - b -> m = -8 + 14 = 6 y = 6x - 14 CHECK -8 = 6 - 14... YES

OpenStudy (mathstudent55):

\(a_{n+1} = a_n + 6\)